40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15133     Entry time: Mon Jan 20 12:16:50 2020
Author: gautam 
Type: Update 
Category: PSL 
Subject: PMC input reverted to AOM zeroth order beam 

Summary:

  1. The input beam to the PMC cavity was changed back to the zeroth order beam from the AOM. 
  2. The PMC was locked and nominal transmission levels were recovered.
  3. The AOM driver voltage was set to 0V DC. 
  4. A razor beam dump was placed to catch the first (and higher order) beams from the AOM (see Attachment #1), but allow the zeroth order beam to reach the PMC cavity.
  5. Some dangling cabling was cleared from the PSL enclosure.

Details

  • HEPA turned to 100% while work was going on in the PSL enclosure.
  • Input power to the PMC cut from ~1.3 W to ~20 mW using the first available HWP downstream of the laser head, before any realignment work was done.
  • Next, the beam dump blocking the undeflected zeroth order beam was removed.
  • Triangle wave was applied to the PZT servo board "EXT DC" input to sweep the cavity length to make the alignment easier.
  • After some patient alignment, I could see a weak transmitted beam locked to some high order mode, at which point I increased the input power to 200mW, and did the fine alignment by looking at the mode shape of the transmitted beam.
  • Once I could lock to a TEM00 mode, I bumped the power back up to the nominal 1.3W, I fine tuned the alignment further by minimizing PMC REFL's DC level. 
  • Dialled the power back down (using HWP) for installation of the beam block to catch the AOM's first (and higher order) beams.
  • Checked that the reflected beam from the PMC cavity is well centered on the PMC REFL PDH photodiode. The ghost from the AR coating of the high-T beamsplitter is blocked by the iris installed by yehonathan on Friday. 
  • The beam was a little low on the PMC REFL CCD camera - I raised the camera by ~1cm.
  • With the beam axis well matched to the PMC, I measured 1.33 mW going into the cavity, and 1.1 W transmitted, so T_{\mathrm{PMC}} \approx 83 \, \%. Whatever loss numbers we extract should be consistent with this fact.
  • HEPA turned back down to 30% shortly after noon.

Note that for all the alignment work, only the two steering mirrors immediately upstream of the PMC cavity were touched.

Attachment 1: IMG_8362.JPG  626 kB  Uploaded Mon Jan 20 15:01:54 2020  | Hide | Hide all
IMG_8362.JPG
ELOG V3.1.3-